# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters and Existence of Optimal Solutions

### 21 Dec 2023 v1

@@ Abstract.
The average cost optimality is known to be a challenging problem for partially observable stochastic control, with few results available beyond the finite state, action, and measurement setup, for which somewhat restrictive conditions are available. In this paper, we present explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. In particular, we present a contraction based analysis, which is new to the literature to our knowledge, building on recent regularity results for non-linear filters.
@@
Non-linear filtering, unique ergodicity, filter stability
@@ 1. Introduction
We let the agent’s goal be to minimize the expected discounted cost of its actions over an infinite period of time.
@@
for some discount factor $\beta\in(0,1)$, over the set of admissible policies $\gamma\in\Gamma$, where $c:\mathbb{X}\times\mathbb{U}\rightarrow\mathbb{R}$ is the stage-wise measurable cost function, and the expectation $E_{\mu}^{\mathcal{T},\gamma}$ is taken over the initial state probability measure $\mu$ and the transition kernel $\mathcal{T}$ under policy $\gamma$. The optimal cost for the discounted infinite horizon is defined as
@@
The average cost control problem under partial observations involves finding the optimal policy that minimizes the average cost of the system over an infinite horizon:
@@
$J(\mu,\gamma)=\limsup_{n\to\infty}\frac{1}{n}E_{\mu}^{\gamma}\left[\sum_{t=0}^{\infty}c\left(X_{t},U_{t}\right)\right].$
@@ 1.1. Notation and preliminaries
It is known that any POMDP can be reduced to a completely observable Markov process [1], [2], whose states are the posterior state distributions or ’beliefs‘ of the observer; that is, the state at time $t$ is
$z_{t}(\,\cdot\,):=P\{X_{t}\in\,\cdot\,|y_{0},\ldots,y_{t},u_{0},\ldots,u_{t-1}\}\in{\mathcal{P}}(\mathbb{X}).$
@@
(5) $\rho_{BL}(\mu,\nu)=\sup_{\|f\|_{BL}\leq 1}\biggl{|}\int_{\mathbb{X}}f(e)\mu(de)-\int_{\mathbb{X}}f(e)\nu(de)\biggr{|},$
@@
$\|f\|_{BL}:=\|f\|_{\infty}+\sup_{e\neq e^{\prime}}\frac{f(e)-f(e^{\prime})}{d_{\mathbb{X}}(e,e^{\prime})},$
and $d_{\mathbb{X}}$ is the metric on $\mathbb{X}$.
When $\mathcal{X}$ is compact, one way to metrize $\mathcal{Z}$ under the weak convergence topology is via the Kantorovich-Rubinstein metric (also known as the Wasserstein metric of order $1$) ([5], Theorem 8.3.2) defined as follows
@@ 1.2. Average Cost Optimality Equation
Average cost optimality equation (ACOE) plays a crucial role for the analysis and the existence results of MDPs under the infinite horizon average cost optimality criteria. The triplet $(h,\rho^{*},\gamma^{*})$, where $h,\gamma:\mathcal{Z}\to\mathds{R}$ are measurable functions and $*\rho*\in\mathds{R}$ is a constant forms the ACOE if
$h(z)+\rho^{*}$ $=\inf_{u\in\mathds{U}}\left\{\tilde{c}(z,u)+\int h(z_{1})\eta(dz_{1}|z,u)\right\}$
@@ 1.3. Statement of the Main Result
    $(\mathbb{X},d)$ is a bounded compact metric space with diameter $D$ (where $D=\sup_{x,y\in\mathbb{X}}d(x,y)$).
@@
As can be seen, the testability/verification of these criteria is not overly demanding. We provide two examples to illustrate this, one in a continuous setting and the other in a discrete setting.
@@ 1.4. Examples
Please note that $\delta(Q)>1/2$, and we can choose $\alpha=1$. Therefore, according to Theorem 1.1, the Average Cost Optimal Equation (ACOE) has an optimal policy.
The second example will be for the continuous state case.
@@
For this example, we have $D=supd(x,x^{\prime})=2$, and by choosing $\alpha=\frac{2}{7}$ and $K_{1}=1$, we can set $K_{2}$ to $6/7$. This choice ensures that we satisfy the conditions outlined in Assumption 1. Consequently, we can directly apply Theorem 1.1 to this specific example.
@@ 1.5. Literature Review and Comparison
In the following we present a detailed literature review and also make a comparison with out results. We will in particular highlight that many of the results in the literature are not easy to verify, which also serves to demonstrate that the problem is a challenging one.
@@
Borkar and Budhiraja’s [13] assumptions and results can be explained as follows. They consider $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{U}$ as Polish spaces, where $\mathbb{X}$ is a finite-dimensional Euclidean space (not necessarily compact, unlike ours) and $\mathbb{U}$ is a compact space. In their setup, unlike ours, the observation measurement depends on the previous time’s state and action. For $(x,u)\in\mathbb{X}\times\mathbb{U}$, the transition probability measure $p(x,u,dz,dy)\in\mathscr{P}(\mathbb{X}\times\mathbb{Y})$ is defined. [13] assumes that $p$ is continuous in a strong sense: Let $\lambda$ denote the Lebesgue measure on $\mathbb{X}$ and let $\eta\in\mathscr{P}(\mathbb{Y})$, assume a density function $\varphi(x,u,z,y)$ on $\mathbb{X}\times\mathbb{U}\times\mathbb{X}\times\mathbb{Y}$ such that $p(x,u,dz,dy)=\varphi(x,u,z,y)\lambda(dz)\eta(dy)$, with $\varphi(\cdot)>0$. It is assumed that $\varphi(x,u,z,y)$ is continuous. Note that by Scheff’e’s lemma, this implies that the transition kernel as well as the measurement kernel are continuous in total variation. The update rule is then given as
@@
There exist $\mathscr{V}\in\mathbb{C}(\mathbb{X})$ and $\hat{\mathscr{V}}\in\mathbb{C}(\mathbb{X})$ satisfying $\lim_{\|x\|\rightarrow\infty}\mathscr{V}(x)=\infty$ and $\lim_{\|x\|\rightarrow\infty}\hat{\mathscr{V}}(x)=\infty$, and under any wide sense admissible $\left\{Z_{n}\right\}$,
$E\left[\hat{\mathscr{V}}\left(X_{n+1}\right)\mid\mathscr{F}_{n}\right]-\hat{\mathscr{V}}\left(X_{n}\right)\leqslant-\mathscr{V}\left(X_{n}\right)+\hat{C}I_{\hat{B}}\left(X_{n}\right),$
$\limsup_{n\rightarrow\infty}\frac{E\left[\hat{\mathscr{V}}\left(X_{n}\right)\right]}{n}=0,$
where $\hat{C}>0$ and $\hat{B}=\{x\in S:\|x\|\leqslant\hat{R}\}$ for some $\hat{R}>0$ and $\mathscr{F}_{n}=\sigma(X_{[0,n]},Y_{[0,n]},U_{[0,n]})$.
@@
$\mathcal{Z}_{\epsilon}=\{\mu\in\mathcal{Z}:\mu(x)>\epsilon\quad\forall x\in X\}$. There exist constants $\epsilon>0$, $k_{0}\in\mathbb{N}$, and $\alpha<1$ such that if $z_{*}(\beta)\in\arg\min_{z\in Z}J_{\beta}(z)$, then for each $\beta\in[\alpha,1)$ we have
@@
There exists $k\geqslant 1$, and $\Delta>0$ such that, for all $y^{k}\in\mathbb{Y}^{k}$ and $u^{k}\in\mathbb{U}^{k}$,
@@
With this review, we have both summarized some existing related work, and also highlighted that our paper presents accessible and testable conditions, compared with most of the literature reviewed above.
@@ 2. Proof of the Main Theorem
Theorem 2.1 ([17], Theorem 7-i).
Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces. If Assumption 1-3 is fulfilled, then we have
$\rho_{BL}\left(\eta(\cdot\mid z),\eta\left(\cdot\mid z^{\prime}\right)\right)\leq 3\left(1+\alpha\right)\rho_{BL}\left(z,z^{\prime}\right).$
Theorem 2.2 ([17], Theorem 7-iv).
$\rho_{BL}\left(\eta(\cdot|z,u),\eta(\cdot|z^{\prime},u)\right)\leq(3-2\delta(O))(1-\tilde{\delta}(\mathcal{T}))\|z-z^{\prime}\|_{TV}$
@@
Building on these results and their proof method, the following result essentially follows from [18]. A proof sketch is given in the appendix.
@@
[18] Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces. If Assumptions 1-1,3 are fulfilled, then we have
@@
[19, Theorem 2] Under Assumption 1-2, the transition probability $\eta(\cdot\mid z,u)$ of the filter process is weakly continuous in $(z,u)$.
We also highlight that additional recent results are available on the weak Feller property, such as those in [20]; however, as there are no restrictions on $Q$, this result [19, Thm.2] is relevant here.
@@
We adopt the approach in the proof of [21, Theorem 4.37]. Let $f\in\operatorname{Lip}(\mathcal{Z},k)$ for some $k>0$. Then $g=\frac{f}{k}\in\operatorname{Lip}(\mathcal{Z},1)$ and therefore for all $u\in\mathrm{U}$ and $z,y\in\mathcal{Z}$ we have
@@
$\leq\max_{u\in\mathrm{U}}\bigg{\{}|c(z,u)-c(y,u)|$
@@
$T^{n}f\in\operatorname{Lip}\left(Z,M_{n}\right),$
where $M_{n}=K_{1}+\beta K_{2}M_{n-1}$ and thus $M_{n}=K_{1}\sum_{i=0}^{n-1}\left(\beta K_{2}\right)^{i}+k\left(\beta K_{2}\right)^{n}$. We have that $M_{n}\to\frac{K_{1}}{1-\beta K_{2}}$ since $K_{2}<1$. Hence, $T^{n}f\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ for all $n$, and therefore, $J_{\beta}^{*}\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ since it is closed with respect to the sup-norm. Taking $k\leq\frac{K_{1}}{1-\beta K_{2}}$, we certify that the fixed point satisfies the desired Lipschitz continuity.
@@
(13) $h_{\beta}(z)=J_{\beta}(z)-J_{\beta}(z_{0})\leq\frac{K_{1}}{1-K_{2}}W_{1}(z,z_{0})\leq\frac{K_{1}\cdot D}{1-K_{2}}.$
In view of the results above, we now introduce a crucial auxiliary theorem that will play a pivotal role in establishing our main result:
@@
    The one stage cost function c is bounded and continuous.
@@
    The stochastic kernel $\mathcal{T}(\cdot\mid x,u)$ is weakly continuous in $(x,u)\in\mathbb{X}\times\mathbb{U}$, i.e., if $\left(x_{k},u_{k}\right)\rightarrow(x,u)$, then $\mathcal{T}\left(\cdot\mid x_{k},u_{k}\right)\rightarrow$ $\mathcal{T}(\cdot\mid x,u)$ weakly.
@@
    $\mathbb{X}$ is $\sigma$-compact, that is, $\mathbb{X}=\cup_{n}S_{n}$ where $S_{n}\subset S_{n+1}$ and each $S_{n}$ is compact.
There exists $\alpha\in(0,1)$ and $N\geq 0$, and a state $z_{0}\in\mathbb{X}$ such that,
@@
    $-N\leq h_{\beta}(z)\leq N$ for all $z\in\mathbb{X}$ and $\beta\in[\alpha,1)$, where
    $h_{\beta}(z)=J_{\beta}(z)-J_{\beta}\left(z_{0}\right),$
    for some fixed $z_{0}\in\mathbb{X}$.
@@
    The sequence $\left\{h_{\beta(k)}\right\}$ is equicontinuous, where $\{\beta(k)\}$ is a sequence of discount factors converging to 1 which satisfies $\lim_{k\rightarrow\infty}(1-\beta(k))J_{\beta(k)}^{*}(z)=\rho^{*}$ for all $z\in\mathbb{X}$ for some $\rho^{*}\in[0,L]$.

### 13 Mar 2024 v2

# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters, Optimal Solutions and Approximations

@@ Abstract.
The average cost optimality is known to be a challenging problem for partially observable stochastic control, with few results available beyond the finite state, action, and measurement setup, for which somewhat restrictive conditions are available. In this paper, we present explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. In particular, we present a new contraction based analysis, which is new to the literature to our knowledge, building on recent regularity results for non-linear filters. Beyond establishing existence, we also present several implications of our analysis that are new to the literature: (i) robustness to incorrect priors (ii) near optimality of policies based on quantized approximations, (iii) near optimality of policies with finite memory, and (iv) convergence in Q-learning. In addition to our main theorem, each of these represents a novel contribution for average cost criteria.
@@
Non-linear filtering, average cost optimality equation
@@ 1. Introduction
We may let the agent’s goal be to minimize the expected discounted cost
@@
for some discount factor $\beta\in(0,1)$, over the set of admissible policies $\gamma\in\Gamma$, where $c:\mathbb{X}\times\mathbb{U}\rightarrow\mathbb{R}$ is the stage-wise measurable cost function, and the expectation $E_{\mu}^{\gamma}$ is taken over the initial state probability measure $\mu$ under policy $\gamma$. The optimal cost for the discounted infinite horizon is defined as
@@
The average cost control problem under partial observations involves finding an optimal policy that minimizes the average cost of the system over an infinite horizon:
@@
$J(\mu,\gamma)=\limsup_{n\to\infty}\frac{1}{n}E_{\mu}^{\gamma}\left[\sum_{t=0}^{n-1}c\left(X_{t},U_{t}\right)\right].$
Contribution:
  • (i)
    In the field of Partially Observable Markov Decision Processes (POMDPs), the average cost optimality equation (ACOE) typically has few findings in the literature, especially in scenarios beyond those with finite states, actions, and measurements, often restricted by stringent conditions, as we discuss in detail in this paper. This paper introduces a novel result concerning the existence of a solution for the ACOE, by utilizing a Wasserstein regularity result (Theorem 2.2), as detailed in Theorem 1.1. Under Assumption 1, we show the existence of a solution for the ACOE. We provide a detailed comparison with the existing results in the literature and highlight the explicit nature of our conditions.
  • (ii)
    Subsequently, the paper then presents several applications and implications of the existence of a solution to ACOE for POMDPs. These represent new contributions in this area of study to the best of our knowledge:
    1. (1)
      Subsection 3.1 establishes robustness to incorrect priors under universal filter stability, demonstrating that an optimal policy designed for an incorrect prior remains optimal when applied for the correct prior under the average cost criteria, leading to a complete robustness property. Additionally, the subsection includes an analysis for the discounted cost setup.
    2. (2)
      Subsection 3.2.1 shows the near-optimality of quantized approximation policies for average cost criteria.
    3. (3)
      Subsection 3.2.2 examines how, under conditions of filter stability, optimal policies derived from a finite window of measurements and actions for the discounted cost criteria are near-optimal solutions for the average cost criteria in POMDPs.
    4. (4)
      We also highlight the use of Q-learning to derive near-optimal policies applicable for both discounted finite window and discounted quantized approximation scenarios.
@@ 1.1. Notation and preliminaries
It is known that any POMDP can be reduced to a completely observable Markov process [1], [2], whose states are the posterior state distributions or ’beliefs‘ of the observer; that is, the state at time $n$ is
$z_{n}(\,\cdot\,):=P\{X_{n}\in\,\cdot\,|y_{0},\ldots,y_{n},u_{0},\ldots,u_{n-1}\}\in{\mathcal{P}}(\mathbb{X}).$
@@
(5) $\rho_{BL}(\mu,\nu)=\sup_{\|f\|_{BL}\leq 1}\biggl{|}\int_{\mathbb{X}}f(x)\mu(dx)-\int_{\mathbb{X}}f(x)\nu(dx)\biggr{|},$
@@
$\|f\|_{BL}:=\|f\|_{\infty}+\left\lVert f\right\rVert_{L},\quad\left\lVert f\right\rVert_{L}=\sup_{x\neq x^{\prime}}\frac{f(x)-f(x^{\prime})}{d_{\mathbb{X}}(x,x^{\prime})}$
and $d$ is the metric on $\mathbb{X}$.
When $\mathbb{X}$ is compact, one way to metrize $\mathcal{Z}$ under the weak convergence topology is via the Kantorovich-Rubinstein metric (also known as the Wasserstein metric of order $1$) ([5], Theorem 8.3.2) defined as follows
@@ 1.2. Average Cost Optimality Equation
Average cost optimality equation (ACOE) plays a crucial role for the analysis and the existence results of MDPs under the infinite horizon average cost optimality criteria. The triplet $(h,\rho^{*},\gamma^{*})$, where $h,\gamma:\mathcal{Z}\to\mathbb{R}$ are measurable functions and $*\rho*\in\mathbb{R}$ is a constant forms the ACOE if
$h(z)+\rho^{*}$ $=\inf_{u\in\mathbb{U}}\left\{\tilde{c}(z,u)+\int h(z_{1})\eta(dz_{1}|z,u)\right\}$
@@ 1.3. Statement of the Main Result
    $U$ is a compact space and $(\mathbb{X},d)$ is a compact metric space with diameter $D$ (where $D=\sup_{x,y\in\mathbb{X}}d(x,y)$).
@@
As can be seen, the testability/verification of these criteria is explicit. We provide three examples to illustrate this, one in a discrete setting and the others in continuous setting.
As we will discuss in more detail in Section 3, the theorem provides significant implications for robustness and approximations. By utilizing this theorem, we can derive near-optimal policies for the average cost criteria in partially observable Markov decision processes. To the best of our knowledge, obtaining near-optimal policies for the average cost criteria is a novel result in the literature on Partially Observable Markov Decision Processes (POMDPs). Additionally, under certain conditions, our findings contribute a new perspective to the literature on robustness, demonstrating that the average cost optimization problem is completely robust to initialization errors.
@@ 1.4. Examples
For this example, please note that $\delta(Q)$ is greater than $1/2$, and the diameter $D$ is equal to 1. We can choose $\alpha$ to be 1. Hence, according to Theorem 1.1, the Average Cost Optimal Equation (ACOE) has a solution.
Alternatively, if we can select $\alpha$ to be less than 1, a more relaxed condition on $\delta(Q)$ would suffice. If we can choose $\alpha$ to be less than $2/3$, without any other constraints on the observation kernel, we can assert that the Average Cost Optimal Equation (ACOE) has a solution. As we will see in the literature review section, these conditions are new.
The second example will be for the continuous state space case.
@@
We have $D=supd(x,x^{\prime})=2$, and by choosing $\alpha=\frac{2}{7}$ and $K_{1}=1$, we can set $K_{2}$ to $6/7$. This choice ensures that we satisfy the conditions outlined in Assumption 1. Consequently, we can directly apply Theorem 1.1 to this specific example.
@@ 1.5. Literature Review and Comparison
In the following we present a detailed literature review and also make a comparison with our results. We will in particular highlight that many of the results in the literature are not easy to verify, which also serves to demonstrate that the problem is a challenging one.
@@
Borkar and Budhiraja’s [13] assumptions and results can be explained as follows. They consider $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{U}$ as Polish spaces, where $\mathbb{X}$ is a finite-dimensional Euclidean space (not necessarily compact, unlike ours) and $\mathbb{U}$ is a compact space. In their setup, unlike ours, the observation measurement depends on the previous time’s state and action. For $(x,u)\in\mathbb{X}\times\mathbb{U}$, the transition probability measure $p(x,u,dz,dy)\in\mathcal{P}(\mathbb{X}\times\mathbb{Y})$ is defined. [13] assumes that $p$ is continuous in a strong sense: Let $\lambda$ denote the Lebesgue measure on $\mathbb{X}$ and let $\eta\in\mathcal{P}(\mathbb{Y})$, assume a density function $\varphi(x,u,z,y)$ on $\mathbb{X}\times\mathbb{U}\times\mathbb{X}\times\mathbb{Y}$ such that $p(x,u,dz,dy)=\varphi(x,u,z,y)\lambda(dz)\eta(dy)$, with $\varphi(\cdot)>0$. It is assumed that $\varphi(x,u,z,y)$ is continuous. Note that by Scheff’e’s lemma, this implies that the transition kernel as well as the measurement kernel are continuous in total variation. The update rule is then given as
@@
There exist $\mathcal{V}\in\mathbb{C}(\mathbb{X})$ and $\hat{\mathcal{V}}\in\mathbb{C}(\mathbb{X})$ satisfying $\lim_{\|x\|\rightarrow\infty}\mathcal{V}(x)=\infty$ and $\lim_{\|x\|\rightarrow\infty}\hat{\mathcal{V}}(x)=\infty$, and under any wide sense admissible $\left\{Z_{n}\right\}$,
$E\left[\hat{\mathcal{V}}\left(X_{n+1}\right)\mid\mathbb{F}_{n}\right]-\hat{\mathcal{V}}\left(X_{n}\right)\leq-\mathcal{V}\left(X_{n}\right)+\hat{C}I_{\hat{B}}\left(X_{n}\right),$
$\limsup_{n\rightarrow\infty}\frac{E\left[\hat{\mathcal{V}}\left(X_{n}\right)\right]}{n}=0,$
where $\hat{C}>0$, $\hat{B}=\{x\in S:\|x\|\leq\hat{R}\}$ for some $\hat{R}>0$ and $\mathbb{F}_{n}=\sigma(X_{[0,n]},Y_{[0,n]},U_{[0,n]})$.
@@
$\mathcal{Z}_{\epsilon}=\{\mu\in\mathcal{Z}:\mu(x)>\epsilon\quad\forall x\in X\}$. There exist constants $\epsilon>0$, $k_{0}\in\mathbb{N}$, and $\alpha<1$ such that if $z_{*}(\beta)\in\arg\min_{z\in\mathcal{Z}}J^{*}_{\beta}(z)$, then for each $\beta\in[\alpha,1)$ we have
@@
There exists $k\geq 1$, and $\Delta>0$ such that, for all $y^{k}\in\mathbb{Y}^{k}$ and $u^{k}\in\mathbb{U}^{k}$,
@@
In [17], the authors study near optimality of finite window policies for average cost problems where the state, action and observation spaces are finite; under the condition that the liminf and limsup of the average cost are equal and independent of the initial state, the paper establishes the near-optimality of (non-stationary) finite memory policies. Here, a concavity argument building on [18] (which becomes consequential by the equality assumption) and the finiteness of the state space is crucial. The paper shows that for any given $\epsilon>0$, there exists an $\epsilon$-optimal finite window policy.
With this review, we have both summarized some existing key studies, and also highlighted that our paper presents accessible and testable conditions, compared with most of the literature reviewed above.
@@ 2. Proof of the Main Theorem
Theorem 2.1 ([19], Theorem 7-i, Theorem 7-iv).
Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces.
  1. i.
    If Assumption 1-3 is fulfilled, then we have
    $\rho_{BL}\left(\eta(\cdot\mid z,u),\eta\left(\cdot\mid z^{\prime},u\right)\right)\leq 3\left(1+\alpha\right)\rho_{BL}\left(z,z^{\prime}\right).$
@@
  2. ii.
    $\rho_{BL}\left(\eta(\cdot|z,u),\eta(\cdot|z^{\prime},u)\right)\leq(3-2\delta(Q))(1-\tilde{\delta}(\mathcal{T}))\|z-z^{\prime}\|_{TV}$
    for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$.
Building on these results and their proof method, the following result follows from [20], Theorem 5.1] which considered the control-free setup. A proof sketch is given in the appendix.
@@
Assume that $\mathbb{X}$ and $\mathbb{Y}$ are Polish spaces. If Assumption 1-1 and Assumption 1-3 are fulfilled, then we have
@@
[21, Theorem 2] Under Assumption 1-2, the transition probability $\eta(\cdot\mid z,u)$ of the filter process is weakly continuous in $(z,u)$.
We also highlight that additional recent results are available on the weak Feller property, such as those in [22]; however, as there are no restrictions on $Q$, this result [21, Thm.2] is relevant here.
@@
We adopt the approach in the proof of [23, Theorem 4.37]. Let $f\in\operatorname{Lip}(\mathcal{Z},k)$ for some $k>0$. Then $g=\frac{f}{k}\in\operatorname{Lip}(\mathcal{Z},1)$ and therefore for all $u\in\mathrm{U}$ and $z,y\in\mathcal{Z}$ we have
@@
$\leq\max_{u\in\mathrm{U}}\bigg{\{}|\tilde{c}(z,u)-\tilde{c}(y,u)|$
@@
$T^{n}f\in\operatorname{Lip}\left(\mathcal{Z},M_{n}\right),$
where $M_{n}=K_{1}+\beta K_{2}M_{n-1}$ and thus $M_{n}=K_{1}\sum_{i=0}^{n-1}\left(\beta K_{2}\right)^{i}+k\left(\beta K_{2}\right)^{n}$. Then, the sequence $M_{n}$ monotonically converges to $\frac{K_{1}}{1-\beta K_{2}}$ since $K_{2}<1$. Hence, $T^{n}f\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ for all $n$, and therefore, $J_{\beta}^{*}\in\operatorname{Lip}\left(\mathcal{Z},\frac{K_{1}}{1-\beta K_{2}}\right)$ since it is closed with respect to the sup-norm. Taking $k\leq\frac{K_{1}}{1-\beta K_{2}}$, we certify that the fixed point satisfies the desired Lipschitz continuity.
@@
(13) $h_{\beta}(z)=J^{*}_{\beta}(z)-J^{*}_{\beta}(z_{0})\leq\frac{K_{1}}{1-K_{2}}W_{1}(z,z_{0})\leq\frac{K_{1}\cdot D}{1-K_{2}}.$
In view of the results above, we now introduce a crucial auxiliary result that will play a pivotal role in establishing our main result:
@@
    The one stage cost function $\tilde{c}$ is bounded and continuous.
@@
    The stochastic kernel $\eta(\cdot\mid x,u)$ is weakly continuous in $(x,u)\in\mathcal{Z}\times\mathbb{U}$, i.e., if $\left(x_{k},u_{k}\right)\rightarrow(x,u)$, then $\eta\left(\cdot\mid x_{k},u_{k}\right)\rightarrow$ $\eta(\cdot\mid x,u)$ weakly.
@@
    $\mathcal{Z}$ is $\sigma$-compact, that is, $\mathcal{Z}=\cup_{n}S_{n}$ where $S_{n}\subset S_{n+1}$ and each $S_{n}$ is compact.
There exists $\alpha\in(0,1)$ and $N\geq 0$, and a state $z_{0}\in\mathcal{Z}$ such that,
@@
    $-N\leq h_{\beta}(z)\leq N$ for all $z\in\mathcal{Z}$ and $\beta\in[\alpha,1)$, where
    (14) $h_{\beta}(z)=J^{*}_{\beta}(z)-J^{*}_{\beta}\left(z_{0}\right).$
@@
    The sequence $\left\{h_{\beta(k)}\right\}$ is equicontinuous, where $\{\beta(k)\}$ is a sequence of discount factors converging to 1 which satisfies $\lim_{k\rightarrow\infty}(1-\beta(k))J_{\beta(k)}^{*}(z)=\rho^{*}$ for all $z\in\mathcal{Z}$ for some $\rho^{*}\in[0,L]$.
Lemma 2.2.
[24, Theorem 7.3.3] Under Assumption 9, a solution to the average cost optimality equation (ACOE) exists, leading to the existence of an optimal control policy, and optimal cost is constant for every initial state.
We note that a similar result, with slightly stronger continuity conditions (not applicable to our setup) under [3, Assumption 4.2.1], can be found in [3, Theorem 5.5.4]. Using this auxiliary result, we are now ready to present the proof of the main theorem.
Proof of Theorem 1.1.
In view of Lemma 2.1, all conditions of Lemma 2.2 are satisfied. Assumption 9-2 follows from Theorem 2.3, and Assumption 9-5 follows from (13). Notably, the equi-continuity condition, Assumption 9-6 holds as a result of the relation
$h_{\beta}(z)=J^{*}_{\beta}(z)-J^{*}_{\beta}(z_{0})\leq\frac{K_{1}}{1-K_{2}}W_{1}(z,z_{0})$
for all $\beta\in(0,1)$. This concludes the proof.
3. Implications for Approximations, Robustness and Learning
In this section, we discuss the implications of our results, namely, that of the existence of a solution to the average cost optimality equation and Wasserstein regularity.
First, we establish robustness; studying how an optimal policy developed for an incorrect initial prior performs when applied to the correct initial distribution in the context of the average cost criteria. Under certain conditions, we find that the average cost optimization problem is robust to errors in initialization.
We then study approximate optimality through the quantization of the belief space, or $P(\mathbb{X})$. This method yields a near-optimal policy for the original problem. We then present an alternate method to construct a finite Markov Decision Process (MDP). This is achieved by replacing the complete observable Markov process with the most recent N observations and actions. Within this finite state MDP framework, we apply Q-learning to obtain the optimal policy.
3.1. Robustness to Incorrect Priors for Discounted and Average Cost Criteria
In this section, we discuss the implications of our results on robustness. Let us first formally define the robustness problem. We have already introduced $J^{*}$ and $J_{\beta}^{*}$ in the first section.
An optimal control policy $\gamma^{\mu}$ for a given prior $\mu$ is a policy that achieves the lowest expected cost over all admissible control policies:
$J\left(\mu,\gamma^{\mu}\right)=\inf_{\gamma\in\Gamma}J(\mu,\gamma)=J^{*}(\mu)$
Consider a scenario where a controller incorrectly assumes the system’s prior to be $\nu$, while in reality, it is $\mu$. In this case, the controller would implement the policy $\gamma^{\nu}$, optimal for $\nu$, but this results in an expected cost of $J\left(\mu,\gamma^{\nu}\right)$. If the controller used the correct policy, the cost could have been $J^{*}(\mu)$. In studying robustness, we are interested in this difference:
$\begin{gathered}J_{\beta}\left(\mu,\gamma^{\nu}_{\beta}\right)-J_{\beta}^{*}(\mu),\\ J\left(\mu,\gamma^{\nu}\right)-J^{*}(\mu)\end{gathered}$
for the discounted cost problem and the average cost problem. Here $\gamma^{\nu}_{\beta}$ is an optimal policy for the initial $\nu$ in the context of discounted cost criteria, while $\gamma^{\nu}$ is an optimal policy for the initial $\nu$ in the context of average cost criteria.
In the literature on POMDPs, there are upper bounds on this error cost for discounted cost problem [25], studies on its relationship with filter stability, and attempts to find uniform upper bounds for average cost problem [26]. There is also work considering how robustness is affected under different transition kernels for discounted cost problems [27]. The following builds on [25, Theorem 3.2].